# DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization

### v1

@@ 4.1 DualContrast disentangles writing style from MNIST handwritten digit images
Figure 4: Qualitative Results of Unsupervised Content-Transformation Disentanglement on LineMod obtained by VITAE[1] (a) and DualContrast (b), respectively.
@@
However, since transformation labels are not present for MNIST, relying entirely on the content prediction performance for disentanglement might be misleading. For example, the degenerate solution achieves even lower $D_{score}(c|s)$ as transformation does not represent anything. To this end, we report qualitative results for baseline methods and our approach DualContrast (Fig. 2). In generating images with varying content and styles, we observe that Harmony, SpatialVAE, and C-VITAE generate many images with erroneous content and styles. On the other hand, DualContrast produces a few minor errors and no extreme mistakes. We also include tSNE embedding of the content factors inferred by the models on the MNIST test dataset associated with class labels (Figure 3). On the embedding space, DualContrast clearly shows superior clustering performance. All of these suggest better content-transformation disentanglement with our approach than existing explicit parameterization approaches.
@@ 4.2 DualContrast Disentangles ViewPoint from Linemod Object Dataset
Figure 5: Qualitative Results of Unsupervised Content-Transformation Disentanglement on Starmen (a) and (b) shows the results obtained with VITAE [1] and DualContrast, respectively.
LineMod is an RGB object recognition and 3D pose estimation dataset [35] with 15 unique object types. It contains 15 unique objects: ‘ape’, ‘bench vise’, ‘bowl’, ‘cam’, ‘can’, ‘cat’, ‘cup’, ‘driller’, ‘duck’, ‘eggbox’, ‘glue’, ‘hole puncher’, ‘iron’, ‘lamp’ and ‘phone’, photographed in a highly cluttered environment. We use a synthetic version of the dataset [38], which has the same objects rendered under different viewpoints. The dataset is publicly available at this url. The dataset has $,31$ 313 images per object category. We used $1,000$ images per category for training and the remaining for testing.
@@
We evaluated the disentanglement scores $D_{\text{score}}$ (Table 1) for DualContrast and baseline approaches, where DualContrast clearly shows the best performance. We used the object identities as content labels during the evaluation.
@@
We also performed qualitative measurement with image transfer and depicted the result of DualContrast and its most relevant (in terms of the generality of transformations) VITAE in Fig 4. We observe that VITAE distorts the images significantly while performing image generation with content-style transfer. On the other hand, DualContrast does not face such issues. Moreover, DualContrast can disentangle different viewpoints of the objects as the transformation factor using only 2D rotation to create contrastive pairs in its model.
4.3 DualContrast disentangles Human Deformation from starmen shapes dataset
Starmen is a 2D synthetic shape dataset of the deformable human body [36] consisting of $10,000$ shape-focused images, which is publicly available at https://zenodo.org/records/5081988 under CC-BY license. Each starmen image is a grayscale image of size $64\times 64$. They are generated by deforming five control points (two on arms, two on legs, one on head) on template human shapes.
We randomly took $6,000$ starmen images to train our models and kept the remaining $4,000$ for testing. We report the qualitative results by VITAE [1] and DualContrast in Figure 5. From the image generation results with the content-transformation transfer, we can observe that DualContrast disentangles the deformation of starmen shapes as transformation and the identity (based on leg thickness) as content, providing a smooth image generation result. On the other hand, VITAE [1] does not perform such meaningful content-transformation disentanglement, generating implausible results.
@@ 4.3 DualContrast disentangles Human Deformation from starmen shapes dataset
4.4 DualContrast Disentangles Protein Identity from states in Cryo-ET Subtomograms
Figure 6: (a) Isosurface visualization of Nucleosome Density Maps. Each density map slightly varies in terms of conformation. They are associated with the PDB IDs in the figure. (b) 3D Slice by slice visualization of several subtomograms of the corresponding nucleosome density maps.
For each protein class, we collected 6 different protein structures from the RCSB PDB website [42]. RCSB PDB is a web server containing the structure of millions of proteins. For nucleosomes, we collected PDB IDs ‘2pyo’, ‘7kbe’, ‘7pex’, ‘7pey’, ‘7xzy’, and ‘7y00’. All of these are different conformations of nucleosomes that slightly vary in composition (Figure LABEL:fig:supp:nuc-dmaps). For sars-cov-2 spike proteins, we collected PDB IDs ‘6vxx’, ‘6vyb’, ‘6xr8’, ‘6xra’, ‘6zox’, and ‘6zp0’. Among them, only ‘6xra’ shows much variation in structure from the other ones, and the rest of the PDB IDs are very similar in structure. For Fatty Acid Synthase (FAS) Unit, we used PDB IDs ‘8prv’, ‘8ps1’, ‘8ps9’, ‘8psj’, ‘8psm’, ‘8psp’. They also vary very slightly in terms of the structure.
After collecting these 18 PDB structures as PDB files, we used EMAN PDB2MRC [43] to create density maps (as MRC file extension) from the PDB files. We create density maps of size $32^{3}$ with $1$ nm resolution. We then randomly rotate and translate each density map and create $1000$ such copies. We then convolve the density maps with CTF with CTF parameters common in experimental datasets (Defocus -5 nm, Spherical Abberation 1.7, Voltage 300 kV). Afterward, we add noise to the convolved density maps so that the SNR is close to $0.1$. Thus, we prepare $18,000$ realistic subtomograms with 3 different protein identities, each with 6 different conformations. We uploaded the entire dataset anonymously at https://zenodo.org/records/11244440 under CC-BY-SA license. 3D slice-by-slice visualizations for sample subtomograms for nucleosomes, spike proteins, and FAS units are provided in Figure 6.
We trained DualContrast and baseline approaches against the protein subtomogram dataset. We could not train VITAE [1] on subtomogram datasets since it did not define any transformation for 3D data. Designing CPAB transformation for 3D data by ourselves was challenging. Among the other baseline approaches, spatialVAE could not distinguish the protein identities with high heterogeneity at all, which is evident by its embedding space UMAP (Figure 7). Only Harmony provided somewhat plausible results apart from DualContrast.
Figure 7: UMAP of content factor learned by SpatialVAE [2] on our subtomogram dataset. The protein identities are indistinguishable from the embedding space.
We observed that the DualContrast content factor perfectly clustered the three protein identities from the dataset (GMM clustering accuracy $0.99$), profoundly disentangling the transformations (Figure 8(b)). We do not see such perfect clustering for Harmony (Figure 8(a)); rather, protein identities with very different compositions get mixed up due to their conformational variation (GMM clustering accuracy $0.70$). We further observe that DualContrast transformation embedding clusters have similar nucleosome conformations, showing disentanglement of conformation from protein identity. On the other hand, in Harmony, the transformation factor only represents rotation and translation by design and can not capture nucleosome conformations. Moreover, even its content embedding for manually selected (not through automatic clustering) nucleosome subtomograms can not show good clustering of conformations (Figure 8(b)).
@@ 4.4 DualContrast Disentangles Protein Identity from states in Cryo-ET Subtomograms
Figure 8: Disentanglement of composition and conformations in cellular subtomogram dataset. (a) UMAP embedding of the content factor of all subtomograms (top) and only nucleosome subtomograms (bottom) with Harmony [4] (b) UMAP embedding of the content factor (top) and transformation factor (bottom) obtained by DualContrast.
Ablation Study: To evaluate the individual contribution of the contrastive losses, we conduct both quantitative and qualitative ablation analyses of DualContrast. We trained (1) DualContrast without any contrastive loss, which is basically a VAE with two latent spaces, (2) DualContrast with only $L=L_{\text{VAE}}+L_{\text{con (z)}}$, and (3) $L=L_{\text{VAE}}+L_{\text{con (c)}}$. We qualitatively and quantitatively evaluated each model. We show the quantitative results in Table 1 and qualitative results of MNIST in Figure 9.
@@
For (1), the $D_{\text{score}}$ is almost the same for both content and transformation factors, indicating equal predictivity of the digit classes by both factors. This is obvious given that the model has no inductive bias to make different factors capture different information. In model (2), using contrastive loss w.r.t. only transformation factor $z$ makes it uninformative of the data. Thus, it provides a small $D(c|z)$ score as desired, but the changing $z$ does not affect the image generation (Figure 9). On the other hand, in the model (3), using contrastive loss w.r.t. only content factor $c$ gives a less informative content factor, a lower $D(c|c)$ score, and a higher $D(c|z)$ score, contrary to what is desired. These results indicate contrastive loss w.r.t to both factors is crucial to get the desired disentanglement.
5 Limitation
Similar to existing content-transformation disentanglement methods [2, 4, 1, 8, 12], DualContrast is applicable to data samples with single objects. This is reasonable, given content and transformation can not be defined for data with multiple objects. Moreover, we limited our study only to shape-focused image datasets, motivated by the problem of conformation disentanglement of protein subtomograms. Consequently, we only used rotation to create contrastive pairs, which worked pretty well in our cases. For image datasets where appearance or illumination plays a crucial role, using only rotation to create contrastive pairs may not work well. However, disentanglement in natural image datasets is a different scenario that is outside of the scope of this paper.
@@ 6 Conclusion
This work focuses on a challenging setting of the unsupervised content-transformation disentanglement problem in shape-focused image datasets, where the transformation latent factor is not explicitly parameterized. To tackle this problem, we propose a novel method termed DualContrast. DualContrast employs generative modeling with a novel contrastive learning strategy that creates positive and negative pairs for content and transformation latent factors. Our extensive experiments show DualContrast’s effective disentanglement of challenging transformation factors across various shape-based image datasets; including simulated cellular subtomograms, where it solved the unexplored problem of isolation of protein conformations from protein identities, as proof of principle.

### v2

@@ 4.1 DualContrast disentangles several writing styles from MNIST images
Figure 3: Qualitative Results of Unsupervised c-z Disentanglement on MNIST obtained by (a) Harmony, (b) SpatialVAE, (c) VITAE, and (d) DualContrast, respectively. Images are generated by the Decoders given content (c) code from the leftmost column images and transformation (z) code from the topmost row images.
Similar to the related works, we start our experiments with the MNIST dataset of handwritten digits. We performed a quantitative comparison between DualContrast and the baseline approaches, as shown in Table 1. We observed that DualContrast achieves higher $D_{score}(c|c)$ compared to other explicitly parameterized approaches [4, 2, 1].
@@
However, since transformation labels are not present for MNIST, relying entirely on the content prediction performance for disentanglement might be misleading. To this end, we report qualitative results for baseline methods and our approach DualContrast (Fig. 3 and Appendix Fig. 10). In generating images with varying c and z codes, we observe that Harmony, SpatialVAE, and C-VITAE generate many images with erroneous content and transformations. On the other hand, DualContrast does not make such mistakes. This again suggests better c-z disentanglement with our approach than existing explicit parameterization approaches.
Moreover, DualContrast disentangles more than in-plane rotation for MNIST. For the digit 00 marked with green box in Fig. 3, Harmony and SpatialVAE simply rotate the image, not capturing the actual writing style of the digit $2$ in the top row. VITAE somewhat represented the transformation with its explicit modeling of piecewise linear transformation. On the other hand, DualContrast properly captured the writing style of above digit $2$.
We included additional qualitative results, including the $\mathbf{c}$ and $\mathbf{z}$ latent space visualization, in the Appendix A.4.
@@ 4.2 DualContrast Disentangles ViewPoint from Linemod Object Dataset
Figure 4: Qualitative Results of Unsupervised c-z Disentanglement on LineMod obtained by (a) Harmony, (b) SpatialVAE, (c) VITAE, and (d) DualContrast, respectively. Images are generated by the Decoders given c code from the leftmost column images and z code from the topmost row images. Additional content-transformation transfer results are available on Appendix Fig. 12.
LineMod is an RGB object recognition and 3D pose estimation dataset [26] with 15 unique object types. The objects are segmented from real-world scenes, and the segmented images visualize the objects from different viewpoints. The entire dataset has $,31$ 313 images per object category. We used $1,000$ images per category for training and the remaining for testing.
@@
We evaluated the scores $D_{\text{score}}$ (Table 1) for DualContrast and baseline approaches, where DualContrast clearly shows the best performance. During evaluation, we used the object identities as ground truth content factors.
@@
We also performed qualitative measurement similar to MNIST and provided the results in Fig. 4 and Appendix 12. We observe that VITAE distorts the images significantly while performing image generation with c-z transfer. On the other hand, DualContrast does not face such issues. SpatialVAE and Harmony only disentangles in-plane rotation given their design; so changing the z code in these methods only rotates the content sample. However, DualContrast can disentangle different viewpoints of the objects as the z code and changing z changes the viewpoint of objects- reflecting the actual transformation present in the dataset.
@@
4.3 DualContrast disentangles protein composition from conformations in Cryo-ET subtomograms and enables their precise identification
We created a realistically simulated [29] cellular cryo-ET subtomogram dataset of 18,000 samples belonging to 3 different protein identities - Nucleosome, Sars-Cov-2 spike protein, and Fatty Acid Synthase (FAS) unit. We collected 6 different conformations or compositional states for each protein identity from RCSB PDB [30]. Among them, only nucleosome composition has subtly visible differences across all the 6 states. For each conformation per protein identity, 1,000 subtomograms were generated. Each subtomogram is of size $32^{3}$ and contains a protein in random orientation and shift with a very low signal-to-noise ratio (Fig. 5). Further details on the dataset properties are provided in the Appendix A.4.
@@
Figure 5: Disentanglement of composition and conformations in cellular subtomogram dataset with slice-by-slice visualization of $4$ sample subtomograms. UMAP embedding of c codes in (a) SpatialVAE, (b) Harmony, and (c) DualContrast. (d) Slice-by-slice visualization of x-y slices in $4$ sample subtomograms. (e) UMAP embedding of c codes in Harmony trained only nucleosome subtomograms. (f) UMAP embedding of z codes in Harmony trained with all subtomograms.
Figure 6: (a,b,c) Structures obtained with RELION refinement for each (4) cluster of subtomograms, whereas the clustering is performed using Gaussian Mixture Modeling (GMM) on $\mathbf{c}$ codes, predicted by (a) SpatialVAE (b) Harmony, and (c) DualContrast. (d, e) Structures obtained with RELION refinement for the nucleosome subtomograms (identified with previous step) using GMM on (d) Harmony content ($\mathbf{c}$) codes, and (e) DualContrast transformation ($\mathbf{z}$) codes.
We trained DualContrast and baseline approaches (except VITAE since it could not be trained without designing a 3D CPAB transformation) against the protein subtomogram dataset. We provided the quantitative results in Table 1 and qualitative results in Fig. 5 and Fig. 6 respectively.
For qualitative results, we first perform UMAP visualization of the latent codes obtained by the models. We observed that the UMAP of c codes in DualContrast perfectly clustered the protein identities from the dataset, profoundly disentangling the transformations (Fig. 5(c)). We do not see such clustering for Harmony (Fig. 5(b)); rather, protein identities with very different compositions get mixed up due to their conformational variation. SpatialVAE latent code UMAP could not cluster the proteins at all (Fig. 5(a)). We further observe that DualContrast z code UMAP clusters have similar nucleosome conformations, showing disentanglement of conformation from protein identity. On the other hand, in Harmony, the transformation factor only represents rotation and translation by design and can not capture nucleosome conformations. Moreover, even its $\mathbf{c}$ code for manually selected (not through automatic clustering) nucleosome subtomograms can not show good clustering of conformations (Fig. 5(b)).
To further demonstrate the application of such disentanglement, we perform downstream subtomogram averaging to identify the structures from latent space clusters. Subtomogram averaging obtains readily identifiable high SNR structures from multiple low SNR subtomograms through iterative alignment and averaging [28]. In our experiments, Gaussian Mixture Models (GMM) were applied to content ($\mathbf{c}$) codes predicted by models, generating clusters of subtomograms that were averaged using RELION [31] (Fig. 6). SpatialVAE failed to identify proteins, while Harmony identified the FAS, nucleosome, and $2$ spike proteins, though $1$ spike protein appeared as an unrealistic mixture of FAS and spike protein. DualContrast successfully identified the FAS, nucleosome, and $2$ distinct spike protein compositions. For nucleosome conformations, GMM clustering of Harmony’s $\mathbf{c}$ codes for nucleosome cluster subtomograms revealed only $1$ nucleosome conformation but mistakenly included a spike protein. In contrast, clustering DualContrast’s $\mathbf{z}$ codes followed by RELION averaging identified $4$ distinct nucleosome conformations with subtle structural changes, showcasing its ability to disentangle protein composition and conformation, which is unattained by the other existing methods.
@@
Ablation Study: To evaluate the individual contribution of the contrastive losses, we conduct both quantitative and qualitative ablation analyses of DualContrast. We trained (1) DualContrast without any contrastive loss, which is basically a VAE with two latent spaces, (2) DualContrast with only $L=L_{\text{VAE}}+L_{\text{con (z)}}$, and (3) $L=L_{\text{VAE}}+L_{\text{con (c)}}$. We qualitatively and quantitatively evaluated each model. We show the quantitative results in Table 1 and qualitative results of MNIST in Fig. 7. We provide qualitative results and a detailed discussion on the ablation in the Appendix A.4.4.
5 Discussions & Limitations
We introduced an unsupervised content-transformation disentanglement method that, for the first time, does not rely on labels or explicit parameterization of transformations. Our method successfully disentangled transformations that cause subtle pixel-space changes, such as variations in writing styles (MNIST), viewpoint changes (LineMod), and, most importantly, conformational changes in proteins from protein-mixture cryo-ET subtomogram datasets. However, it is not guaranteed to disentangle all transformations, particularly those causing large changes in the pixel space. Instead, the method may classify transformations causing large pixel-space changes as content. Nonetheless, this aligns with scientific imaging, where large changes often reflect compositional shifts and subtle changes represent conformational variations. Disentangling transformations with large pixel-space changes in the pixel space without any explicit parameterization is extremely challenging and might not be practically achievable, but it remains an avenue for future research.
@@ 6 Conclusion
This work focuses on a challenging setting of the unsupervised content-transformation disentanglement problem in scientifically important shape focused image datasets, where the transformation latent code is not explicitly parameterized. To tackle this problem, we propose a novel method termed DualContrast. DualContrast employs generative modeling with a novel contrastive learning strategy that creates positive and negative pairs for content and transformation latent codes. Our extensive experiments show DualContrast’s effective disentanglement of challenging transformations across various shape-based image datasets; including simulated cellular subtomograms, where it solved the unexplored problem of isolation of protein conformations from protein identities, as proof of principle.